# FedMentor: Domain-Aware Differential Privacy for Heterogeneous Federated LLMs in Mental Health

### v1

@@ Abstract
Privacy-preserving adaptation of Large Language Models (LLMs) in sensitive domains (e.g., mental health) requires balancing strict confidentiality with model utility and safety. We propose FedMentor, a federated fine-tuning framework that integrates Low-Rank Adaptation (LoRA) and domain-aware Differential Privacy (DP) to meet per-domain privacy budgets while maintaining performance. Each client (domain) applies a custom DP noise scale proportional to its data sensitivity, and the server adaptively reduces noise when utility falls below a threshold. In experiments on three mental health datasets, we show that FedMentor improves safety over standard Federated Learning without privacy, raising safe output rates by up to three points and lowering toxicity, while maintaining utility (BERTScore F1 and ROUGE-L) within 0.5% of the non-private baseline and close to the centralized upper bound. The framework scales to backbones with up to 1.7B parameters on single-GPU clients, requiring <173<173 MB of communication per round. FedMentor demonstrates a practical approach to privately fine-tune LLMs for safer deployments in healthcare and other sensitive fields.
@@ 1 Introduction
Mental health arises from interacting cognitive, affective, and behavioral processes that shape individual functioning and societal stability. Demand for scalable support has accelerated interest in LLMs for conversational assistance Yang et al. (2023a); Xu et al. (2024); Yang et al. (2024). Deployment remains challenging due to strict privacy requirements, limited interpretability, and legal constraints under HIPAA and GDPR May and Denecke (2022); Nicholas et al. (2020); United States Congress (1996); European Parliament and Council of the European Union (2016). User inputs may include explicit self harm ideation or other clinical signals that require strong confidentiality (e.g., [self-harm ideation example]). These concerns make it difficult to use traditional centralized training methods on sensitive user data, highlighting the need for privacy-preserving and communication-efficient techniques for adapting LLMs in healthcare settings.
@@
To investigate privacy and communication challenges in adapting LLMs to sensitive mental health domains, we present FedMentor, a federated framework that assigns domain-aware Differential Privacy budgets and fine-tunes only LoRA adapters, aggregated with FedAvg McMahan et al. (2017), to achieve private adaptation while preserving utility and fairness under heterogeneity. The overall architecture is shown in Figure 1. FedMentor is designed for scenarios that demand distinct privacy guarantees across domains and where non-IID data can introduce drift and unequal outcomes. We evaluate the framework on three datasets (Dreaddit Turcan and McKeown (2019), IRF Garg et al. (2023), MultiWD Garg et al. (2024)) and five backbones (MobileLLM-ParetoQ-350M Liu et al. (2024a), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025)). FedMentor improves safety with minimal loss in utility: for instance, on MultiWD the Toxicity Safe Rate rises by two points while toxicity decreases from 2.92 to 1.98 (Table 1), and BERTScore F1 Zhang et al. (2020) and ROUGE-L Lin (2004) remain close to standard FL and near the centralized upper bound. Client-level relevance remains consistent across domains, and efficiency is achieved by communicating only adapter weights through LoRA, keeping computation and bandwidth feasible for single-GPU clients. To the best of our knowledge, this is the first study that combines domain-aware Differential Privacy Dwork et al. (2014); Abadi et al. (2016) with federated LoRA fine-tuning for large language models in mental health.
@@
  • Domain-aware private adaptation for mental health. We formulate a federated LoRA approach with per-domain $(\epsilon,\delta)$ that protects sensitive mental health text while preserving task utility. FedMentor raises safety (e.g., +2 percentage points Toxicity Safe Rate with lower toxicity) and keeps BERTScore F1 and ROUGE-L close to FL and near centralized training.
@@ 2 FedMentor foundations: problem and notation
$\Delta W_{\text{LoRA}}^{*}=\arg\min_{\Delta W_{\text{LoRA}}}\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}\,\mathcal{L}_{k}(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}};\mathcal{D}_{k}),$ (1)
@@
$\tilde{\Delta W}_{\text{LoRA}}^{k}=\Delta W_{\text{LoRA}}^{k}+\mathcal{N}\left(0,\left(\frac{\sigma_{l}}{\varepsilon_{d_{k}}}\right)^{2}I\right),\quad\text{satisfying }(\varepsilon_{d_{k}},\delta)\text{-DP}$ (2)
@@
$\mathcal{U}_{m}(\mathcal{M}_{\text{global}})\geq\tau_{m},\quad\forall m\in\{\text{BERTScore},\text{SafeRate},\text{Relevance},\text{Perplexity}\}$ (3)
@@
$P_{d_{i}}(x,y)\neq P_{d_{j}}(x,y),\quad\forall i\neq j,\;\;d_{i},d_{j}\in\{\text{IRF},\text{Dreaddit},\text{MultiWD}\}.$ (4)
@@
$\Delta W=BA,\quad B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k},\quad r\ll\min(d,k)$ (5)
@@ 3.1 Client training
$\Delta W_{\text{LoRA}}^{k}\leftarrow\Delta W_{\text{LoRA}}^{k}-\eta\cdot\nabla_{\Delta W}\mathcal{L}(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}}^{k};b)$ (6)
@@ 3.2 Server aggregation
The server aggregates noised client updates $\{\tilde{\Delta W}_{\text{LoRA}}^{k}\}_{k=1}^{K}$ using dataset-weighted FedAvg McMahan et al. (2017):
$\Delta W_{\text{LoRA}}^{(r)}=\sum_{k=1}^{K}\alpha_{k}\,\tilde{\Delta W}_{\text{LoRA}}^{k},\quad\alpha_{k}=\frac{|\mathcal{D}_{k}^{\text{train}}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}^{\text{train}}|}$ (10)
@@ 3.3 Communication efficiency
6:  Broadcast LoRA weights $\Delta W_{\text{LoRA}}^{(r-1)}$ to all clients $\triangleright$ Server-Update
7:  for client $k\in\{1,\ldots,K\}$ in parallel do $\triangleright$ Client-Update
8:   Attach received $\Delta W_{\text{LoRA}}^{(r-1)}$ to frozen $\Theta_{\text{frozen}}$
9:   for epoch $e=1$ to $E$ do $\triangleright$ Train LoRA only
10:     for batch $b\subset\mathcal{D}_{k}^{\text{train}}$ do
11:      Update: $\Delta W_{\text{LoRA}}^{k}\leftarrow\Delta W_{\text{LoRA}}^{k}-\eta\nabla_{\Delta W}\mathcal{L}(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}}^{k};b)$
12:     end for
13:   end for
14:   $\varepsilon_{k}\coloneqq\mathcal{B}_{d_{k}}$ $\triangleright$ Get domain privacy budget
15:   Add noise: $\tilde{\Delta W}_{\text{LoRA}}^{k}\leftarrow\Delta W_{\text{LoRA}}^{k}+\mathcal{N}(0,(\sigma_{l}/\varepsilon_{k})^{2}I)$
16:   Send noised LoRA weights $\tilde{\Delta W}_{\text{LoRA}}^{k}$ to server
@@
18:  Collect $\{\tilde{\Delta W}_{\text{LoRA}}^{k}\}_{k=1}^{K}$ from clients $\triangleright$ Client Aggregation
19:  $\Delta W_{\text{LoRA}}^{(r)}\leftarrow\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}^{\text{train}}|}{\sum_{j}|\mathcal{D}_{j}^{\text{train}}|}\tilde{\Delta W}_{\text{LoRA}}^{k}$ $\triangleright$ FedAvg
20:  If $\mathcal{U}_{m}<\tau_{m}$ for any $m$: $\sigma_{l}\leftarrow 0.8\cdot\sigma_{l}$ $\triangleright$ Adapt noise
21:  Update budget: $\mathcal{B}_{d}\leftarrow\mathcal{B}_{d}-0.1\varepsilon_{d}$ for each domain $d$
@@
23:return $\mathcal{M}^{(R)}$ with final LoRA weights $\Delta W_{\text{LoRA}}^{(R)}$
@@ 4 Experiment settings
Datasets. We evaluate FedMentor on three mental health datasets: Dreaddit Turcan and McKeown (2019) for stress detection, Interpersonal Risk Factors (IRF) Garg et al. (2023) for Thwarted Belongingness (TBe) and Perceived Burdensomeness (PBu), and MultiWD Garg et al. (2024) for multi-label wellness prediction. Together, these corpora cover complementary tasks and form a natural non-IID benchmark across domains. Regulatory requirements under HIPAA and GDPR restrict centralized collection, and lengthy institutional approvals further limit access, producing small, fragmented datasets Act et al. (1996); Regulation (2018); Nicholas et al. (2020). Prior FL work mitigates performance degradation from heterogeneity with methods such as SCAFFOLD Karimireddy et al. (2020), FedNova Wang et al. (2020), and Adaptive FedOpt Reddi et al. (2021). We adopt a domain-aware FL setting in which each dataset defines a client domain and use this setting to evaluate FedMentor under heterogeneous conditions. Detailed dataset statistics and preprocessing appear in Appendix B.1.
@@
Models. We fine-tune five lightweight language models under federated settings, focusing on compact architectures suited for resource-constrained environments. The models range from 350M to 1.7B parameters: MobileLLM-ParetoQ-350M Liu et al. (2024a), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025). These backbones integrate quantization and distillation to enhance efficiency, enabling deployment on edge devices and GPUs with limited memory. Their compact size further supports faster client-level fine-tuning and inference, which is critical under the resource and communication constraints of FL Ro et al. (2022); Lin et al. (2022); Wang et al. (2024a). To ensure fairness, FedMentor and all baselines are evaluated on identical model architectures so that observed differences arise solely from training strategies. Additional architectural details are provided in Appendix B.2.
Baselines. We compare three setups under identical backbones and LoRA ranks (here, w/o denotes without). (i) Centralized (w/o FL, w/o DP): pool all datasets and fine-tune a single LoRA adapter on the combined corpus, providing an optimistic upper bound. (ii) Federated with LoRA (w/o DP): run FedAvg with frozen backbones and client-specific LoRA adapters; the server aggregates adapter parameters using averaging weighted by data size to isolate decentralization effects. (iii) Federated LoRA under domain-aware DP (FedMentor): clients allocate domain-specific privacy budgets with layer- and adapter-specific noise scaling; the server performs FedAvg over noise-added adapters and reduces noise when utility proxies fall below thresholds. Full baseline configurations appear in Appendix B.3.
Evaluation metrics. We evaluate centralized models with toxicity mean and max, safe rate Gehman et al. (2020), BERTScore F1 Zhang et al. (2020), relevance Reimers and Gurevych (2019), ROUGE-L Lin (2004), and perplexity Jelinek et al. (1977). In federated settings we report toxicity mean, safe rate, BERTScore F1, relevance, and ROUGE-L. With DP we additionally track domain sensitivity, training time, final losses, memory, adapter size, and round-level averages including communication overhead measured by LoRA updates. Appendix B.4 provides a detailed account of the evaluation metrics.
Prompt construction. We use two prompting protocols to ensure consistent evaluation across datasets and models. Zero-shot evaluation Kojima et al. (2022) tests instruction following without the use of in-context examples, relying only on task instructions. Building on this, few-shot evaluation Brown et al. (2020); Sanh et al. (2022) performs supervised LoRA adaptation on each client under domain-aware Differential Privacy with FedAvg aggregation, and inference reuses the same wrapper employed in the zero-shot setting. Full templates, wrappers, and examples appear in Appendix B.5.
Table 2: Efficiency comparison of LLM backbones under FedMentor. Columns report adapter size (Adpt), communication per round (Comm), peak GPU memory (Mem), and training time per round (Time).
@@
Figure 2: Cross-dataset consistency (FS, ROUGE-L) under FL (top) and FedMentor (bottom).

### v2

@@ Abstract
Privacy-preserving adaptation of Large Language Models (LLMs) in sensitive domains (e.g., mental health) requires balancing strict confidentiality with model utility and safety. We propose FedMentor, a federated fine-tuning framework that integrates Low-Rank Adaptation (LoRA) and domain-aware Differential Privacy (DP) to meet per-domain privacy budgets while maintaining performance. Each client (domain) applies a custom DP noise scale proportional to its data sensitivity, and the server adaptively reduces noise when utility falls below a threshold. In experiments on three mental health datasets, we show that FedMentor improves safety over standard Federated Learning (FL) without privacy, raising safe output rates by up to three points and lowering toxicity, while maintaining utility (BERTScore F1 and ROUGE-L) within 0.5% of the non-private baseline and close to the centralized upper bound. The framework scales to backbones with up to 1.7B parameters on single-GPU clients, requiring <173<173 MB of communication per-round. FedMentor demonstrates a practical approach to privately fine-tune LLMs for safer deployments in healthcare and other sensitive fields.
FedMentor: Domain-Aware Differential Privacy for Heterogeneous Federated LLMs in Mental Health
Nobin Sarwar, Shubhashis Roy Dipta University of Maryland, Baltimore County {sms2, sroydip1}@umbc.edu
@@ 1 Introduction
Mental health arises from interacting cognitive, affective, and behavioral processes that shape individual functioning and societal stability. Demand for scalable support has accelerated interest in LLMs for conversational assistance Yang et al. (2023a); Xu et al. (2024); Yang et al. (2024). Deployment remains challenging due to strict privacy requirements, limited interpretability, and legal constraints under HIPAA and GDPR May and Denecke (2022); Nicholas et al. (2020); United States Congress (1996); European Parliament and Council of the European Union (2016). User inputs may include explicit self harm ideation or other clinical signals that require strong confidentiality (e.g., [self-harm ideation example]). These constraints make traditional centralized training challenging for sensitive user data, highlighting the need for techniques that adapt models in a privacy-preserving and communication-efficient manner for healthcare settings. As shown in Figure 1, our approach addresses these needs by combining domain-aware privacy, robustness under non-IID client distributions, and parameter-efficient updates.
Figure 1: Comparison between traditional LLM-based methods for mental health and FedMentor. The baseline lacks privacy protection, robustness to heterogeneity, and communication efficiency, whereas FedMentor introduces domain-aware privacy, achieves robustness under non-IID data, and improves efficiency through LoRA-only updates.
@@
To investigate privacy and communication challenges in adapting LLMs to sensitive mental health domains, we present FedMentor, a federated framework that assigns domain-aware Differential Privacy budgets and fine-tunes only LoRA adapters, aggregated with FedAvg McMahan et al. (2017), to achieve private adaptation while preserving utility and fairness under heterogeneity. The overall architecture is shown in Figure 2. FedMentor is designed for scenarios that demand distinct privacy guarantees across domains and where non-IID data can introduce drift and unequal outcomes. We evaluate the framework on three datasets (Dreaddit Turcan and McKeown (2019), IRF Garg et al. (2023), MultiWD Garg et al. (2024)) and five backbones (MobileLLM-ParetoQ-350M Liu et al. (2024b), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025)). FedMentor improves safety with minimal loss in utility: for instance, on MultiWD the Toxicity Safe Rate (TSR) rises by two points while toxicity decreases from 2.92 to 1.98 (Table 1), and BERTScore F1 Zhang et al. (2020) and ROUGE-L Lin (2004) remain close to standard FL and near the centralized upper bound. Client-level relevance remains consistent across domains, and efficiency is achieved by communicating only adapter weights through LoRA, keeping computation and bandwidth feasible for single-GPU clients. To the best of our knowledge, this is the first study that combines domain-aware DP Dwork et al. (2014); Abadi et al. (2016) with federated LoRA fine-tuning for large language models in mental health.
@@
  • Domain-aware private adaptation for mental health. We formulate a federated LoRA approach with per-domain $(\epsilon,\delta)$ that protects sensitive mental health text while preserving task utility. FedMentor raises safety (e.g., +2 percentage points TSR with lower toxicity) and keeps BERTScore F1 and ROUGE-L close to FL and near centralized training.
@@ 2 FedMentor Foundations: Problem and Notation
$\Delta W_{\text{LoRA}}^{*}$ $=\arg\min_{\Delta W_{\text{LoRA}}}$ (1)
$\;\;\sum_{k=1}^{K}\tfrac{|\mathcal{D}_{k}|}{|\mathcal{D}|}\,\mathcal{L}_{k}\!\left(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}};\,\mathcal{D}_{k}\right),$
@@
$\tilde{\Delta W}_{\text{LoRA}}^{k}$ $=\Delta W_{\text{LoRA}}^{k}+\mathcal{N}\!\left(0,\left(\frac{\sigma_{l}}{\varepsilon_{d_{k}}}\right)^{2}I\right),$ (2)
$\qquad\text{satisfying }(\varepsilon_{d_{k}},\delta)\text{-DP},$
@@
$\mathcal{U}_{m}(\mathcal{M}_{\text{global}})$ $\geq\tau_{m},$ (3)
$\forall m$ $\in\bigl\{\text{BERTScore},\text{SafeRate},$
$\quad\text{Relevance},\text{Perplexity}\bigr\},$
@@
$P_{d_{i}}(x,y)$ $\neq P_{d_{j}}(x,y),\quad\forall i\neq j,$ (4)
$d_{i},d_{j}$ $\in\{\text{IRF},\text{Dreaddit},\text{MultiWD}\}.$
@@
$\Delta W$ $=BA,$ (5)
$B$ $\in\mathbb{R}^{d\times r},\quad A\in\mathbb{R}^{r\times k},\quad r\ll\min\{d,k\}.$
@@ 3.1 Client Training
$\Delta W_{\text{LoRA}}^{k}$ $\leftarrow\Delta W_{\text{LoRA}}^{k}$ (6)
$\hskip-5.0pt-\eta\cdot\nabla_{\Delta W}\,\mathcal{L}\!\left(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}}^{k};\,b\right)$
@@ 3.2 Server Aggregation
The server aggregates noised client updates using dataset-weighted FedAvg McMahan et al. (2017):
$\{\tilde{\Delta W}_{\text{LoRA}}^{k}\}_{k=1}^{K}.$ (10)
$\Delta W_{\text{LoRA}}^{(r)}$ $=\sum_{k=1}^{K}\alpha_{k}\,\tilde{\Delta W}_{\text{LoRA}}^{k},$ (11)
$\alpha_{k}$ $=\frac{|\mathcal{D}_{k}^{\text{train}}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}^{\text{train}}|}$
@@ 3.3 Communication Efficiency
6:   Broadcast LoRA weights $\Delta W_{\text{LoRA}}^{(r-1)}$ to all clients
[…]
8:  for client $k\in\{1,\ldots,K\}$ in parallel do
[…]
10:   Attach $\Delta W_{\text{LoRA}}^{(r-1)}$ to frozen $\Theta_{\text{frozen}}$
11:   for epoch $e=1$ to $E$ do
12:     for batch $b\subset\mathcal{D}_{k}^{\text{train}}$ do
13:      Update:
14:       $\Delta W_{\text{LoRA}}^{k}\leftarrow\Delta W_{\text{LoRA}}^{k}$
15:       $\eta\nabla_{\Delta W}\,\mathcal{L}\!\big(\Theta_{\text{frozen}}+\Delta W_{\text{LoRA}}^{k};\,b\big)$
16:     end for
@@
18:   $\varepsilon_{k}\coloneqq\mathcal{B}_{d_{k}}$
19:   Add noise:
20:     $\tilde{\Delta W}_{\text{LoRA}}^{k}\leftarrow\Delta W_{\text{LoRA}}^{k}+\mathcal{N}\!\big(0,(\sigma_{l}/\varepsilon_{k})^{2}I\big)$
21:   Send noised LoRA weights $\tilde{\Delta W}_{\text{LoRA}}^{k}$ to server
@@
23:   Collect $\{\tilde{\Delta W}_{\text{LoRA}}^{k}\}_{k=1}^{K}$ from clients
[…]
25:  $\Delta W_{\text{LoRA}}^{(r)}\leftarrow\sum_{k=1}^{K}\dfrac{|\mathcal{D}_{k}^{\text{train}}|}{\sum_{j}|\mathcal{D}_{j}^{\text{train}}|}\;\tilde{\Delta W}_{\text{LoRA}}^{k}$
26:  If $\mathcal{U}_{m}<\tau_{m}$ for any $m$:  $\sigma_{l}\leftarrow 0.8\,\sigma_{l}$
27:  Update budget:
28:     $\mathcal{B}_{d}\leftarrow\mathcal{B}_{d}-0.1\,\varepsilon_{d}$ for each domain $d$
29:end for
30:return $\mathcal{M}^{(R)}$ with final LoRA weights $\Delta W_{\text{LoRA}}^{(R)}$
@@ 4 Experiment Settings
Datasets. We evaluate FedMentor on three mental health datasets: Dreaddit Turcan and McKeown (2019) for stress detection, Interpersonal Risk Factors (IRF) Garg et al. (2023) for Thwarted Belongingness (TBe) and Perceived Burdensomeness (PBu), and MultiWD Garg et al. (2024) for multi-label wellness prediction. Together, these corpora cover complementary tasks and form a natural non-IID benchmark across domains. Regulatory requirements under HIPAA and GDPR restrict centralized collection, and lengthy institutional approvals further limit access, producing small, fragmented datasets Act et al. (1996); Regulation (2018); Nicholas et al. (2020). Prior FL work mitigates performance degradation from heterogeneity with methods such as SCAFFOLD Karimireddy et al. (2020), FedNova Wang et al. (2020), and Adaptive FedOpt Reddi et al. (2021). We adopt a domain-aware FL setting where each dataset defines a client domain and use this setting to evaluate FedMentor under heterogeneous conditions. App. B.1 provides dataset statistics and preprocessing details.
Models. FL places tight limits on client compute, memory, and upload bandwidth; therefore, we adopt compact LLM backbones. We fine-tune five lightweight models from 350M to 1.7B parameters: MobileLLM-ParetoQ-350M Liu et al. (2024b), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025). These backbones integrate quantization and distillation to enhance efficiency, enabling deployment on edge devices and GPUs with limited memory. Their compact size further supports faster client-level fine-tuning and inference, which is critical under the resource and communication constraints of FL Ro et al. (2022); Lin et al. (2022); Wang et al. (2024b). To ensure fairness, FedMentor and all baselines are evaluated on identical model architectures so that observed differences arise solely from training strategies. Additional architectural details are provided in App. B.2.
Baselines. We compare three setups under identical backbones and LoRA ranks (here, w/o denotes without). (i) Centralized (w/o FL, w/o DP): pool all datasets and fine-tune a single LoRA adapter on the combined corpus, providing an optimistic upper bound. (ii) Federated with LoRA (w/o DP): run FedAvg with frozen backbones and client-specific LoRA adapters; the server aggregates adapter parameters using data-size weighting to isolate decentralization effects. (iii) Federated LoRA under domain-aware DP (FedMentor): clients allocate domain-specific privacy budgets with layer- and adapter-specific noise scaling; the server performs FedAvg on noised adapters and reduces noise when utility proxies fall below thresholds. Full baseline configurations appear in App. B.3.
@@
Evaluation Metrics. We evaluate centralized models with toxicity mean and max, safe rate Gehman et al. (2020), BERTScore F1 Zhang et al. (2020), relevance Reimers and Gurevych (2019), ROUGE-L Lin (2004), and perplexity Jelinek et al. (1977). In federated settings we report toxicity mean, safe rate, BERTScore F1, relevance, and ROUGE-L. With DP we additionally track domain sensitivity, training time, final losses, memory, adapter size, and round-level averages including communication overhead measured by LoRA updates. App. B.4 provides details on the evaluation metrics.
(a) FL
(b) FedMentor
Figure 3: Cross-dataset consistency (FS, ROUGE-L) under FL (left) and FedMentor (right).
Prompt Construction. We use two prompting protocols to ensure consistent evaluation across datasets and models. Zero-shot evaluation Kojima et al. (2022) tests instruction following without the use of in-context examples, relying only on task instructions. Building on this, few-shot evaluation Brown et al. (2020); Sanh et al. (2022) performs supervised LoRA adaptation on each client under domain-aware DP with FedAvg aggregation, and inference reuses the same wrapper employed in the zero-shot setting. Full templates, wrappers, and examples appear in App. B.5.
@@
Table 2: Efficiency comparison of LLM backbones under FedMentor. Columns report adapter size (Adpt), communication per-round (Comm), peak GPU memory (Mem), and training time per-round (Time).